# The Price equation reveals a universal force-metric-bias law of algorithmic learning and natural selection

### v2

@@
Steven A. Frank**footnotemark: *Diverse learning algorithms, optimization methods, and natural selection share a common mathematical structure, despite their apparent differences. Here I show that a simple notational partitioning of change by the Price equation reveals a universal force-metric-bias (FMB) law: $\Delta\bm{\theta}=\bm{\mathrm{M}}\,\bm{\mathrm{f}}+\bm{\mathrm{b}}+\bm{\xi}$. The force $\bm{\mathrm{f}}$ drives improvement in parameters, $\Delta\bm{\theta}$, in proportion to the slope of performance with respect to the parameters. The metric $\bm{\mathrm{M}}$ rescales movement by inverse curvature. The bias $\bm{\mathrm{b}}$ adds momentum or changes in the frame of reference. The noise $\bm{\xi}$ enables exploration. This framework unifies natural selection, Bayesian updating, Newton’s method, stochastic gradient descent, stochastic Langevin dynamics, Adam optimization, and most other algorithms as special cases of the same underlying process. The Price equation also reveals why Fisher information, Kullback-Leibler divergence, and d’Alembert’s principle arise naturally in learning dynamics. By exposing this common structure, the FMB law provides a principled foundation for understanding, comparing, and designing learning algorithms across disciplines.
@@ Introduction
Learning algorithms pervade modern science. Machine learning improves neural networks through gradient descent. Evolution improves organisms through natural selection. Bayesian inference improves beliefs through probability updates. Despite decades of research in each field, the ultimate relations between these approaches remains unclear.
@@
These connections reflect a deeper geometric principle. Any learning algorithm faces the same fundamental challenge: maximize improvement in performance minus a cost paid for distance moved in the parameter space. Here, movement must be measured appropriately in the parameter space, which is often curved. Curvature may arise from constraints on movement, for example, when a biological population lacks genetic variation in a particular direction. Or curvature may arise from the nonlinear relation between parameters and performance.
@@
The geometric structure of learning has been partially recognized in prior work. Fisher 9 and Rao 10 established that statistical parameter spaces have intrinsic curvature. Amari 11 used this insight to develop natural gradient methods.
@@
This section’s alternative perspectives of dynamics adds a complementary axis to the local-to-population spectrum of methods introduced earlier. I develop this dynamics axis at the population scale, providing the most general conceptual frame. That population context prepares the ground for later discussion of particular algorithms, many of which blend local geometry with broader spatial or temporal scope.
@@
Equivalently, among all possible frequency changes, $\Delta\bm{\mathrm{q}}$, that produce the same improvement in performance, $\Delta_{\bm{\mathrm{f}}}\mskip 2.0mu\bar{w}$, the actual frequency changes for the given trait covariance matrix, $\bm{\mathrm{M}}$, lead to the shortest Fisher-Rao length. In biology, these optimality results are trait-based analogs of Fisher’s Fundamental Theorem for genetics 30, 31, 32.
Of course, forces other than intrinsic performance can alter frequencies. The more we know about those other forces, such as environmental shifts or directional mutation in biology, the more accurately we can account for the consequences of those forces and improve inductive inference about the causal relation between parameters and performance 33, 34, 35.
@@
A particular algorithm can, of course, choose to modify how components of an update are calculated. However, populations set the foundation for analysis. Changes in parameter mean values, $\Delta\bar{\bm{\theta}}$, summarize updates. The metric $\bm{\mathrm{M}}$ typically arises from the covariance of alternative parameter values. Some methods use variational optimization and analogies with free energy, which links learning to various physical principles of dynamics 36.
Second, many algorithms update a single parameter vector. These methods fill in the other end of the spectrum and its middle ground. The purely local strategy, such as Newton’s method, uses a metric and force gradient analyzed at a single point. Hybrid strategies choose metrics that incorporate a broader spatial scope, such as trust regions 37, 38 or the natural gradient 11.
@@
Finally, modern optimizers often include temporal scope. Extensions to stochastic gradient descent, such as Adam 8 and Nesterov 39, 40, explicitly use the history of updates to calculate a bias term, $\bm{\mathrm{b}}$. In these cases, bias explicitly applies physical notions of momentum, in which past movement of parameter values can push future updates beyond local traps on the performance surface.
@@
Covariance matrix algorithms often provide a good black-box optimization method 41. These evolution strategy (ES) methods proceed by analogy with biological evolution. One starts with a target parameter vector and then samples the performance values for a set of different parameter combinations around the target. That set forms a population from which one can calculate an updated target parameter by an empirically calculated covariance matrix and performance gradient, as in eqn 6.
These methods have three challenges: choosing sample points around the current target, estimating the performance gradient, and calculating the new target from the covariance matrix and performance gradient. I briefly summarize how the popular CMA-ES method handles these three challenges 42.
@@
The same $\bm{\mathrm{M}}\,\bm{\mathrm{f}}$ structure occurs in other evolution strategy (ES) learning algorithms, which include natural evolution strategies 43, large-scale parallel ES 44, and separable low-rank CMA-ES 45. The next section moves to the local end of the spectrum, showing that the $\bm{\mathrm{M}}\,\bm{\mathrm{f}}$ structure remains when the population collapses to a single vector.
@@
The Broyden–Fletcher–Goldfarb–Shanno (BFGS) algorithm and its variants are widely used 46. After each step the method compares how the gradient changed to how the parameters moved. That comparison provides information about curvature.
@@
Newton’s local calculation of the curvature via the inverse Hessian is fragile. The local Hessian may not exist, it may change significantly over space, or it may not be invertible to provide the inverse curvature. Trust region methods may solve some of these problems by defining a local region of analysis 37, 38.
@@
The previous examples in this section match the $\bm{\mathrm{M}}\,\bm{\mathrm{f}}$ force-metric pattern in a simple and direct way. By contrast, the mirror descent algorithm has a more complicated geometry that does not obviously map directly to our $\bm{\mathrm{M}}\,\bm{\mathrm{f}}$ pattern 47, 48. However, the following derivations show that this algorithm does in fact use the same basic force-metric approach but with a more involved method to obtain the curvature metric.
@@
This update is a first-order approximation of the mirror descent method 47, 48. In effect, we are making a Newton-like step in an alternative mirror geometry with metric $\bm{\mathrm{H}}^{-1}_{\varphi}$, then mapping that step back to our original geometry for $\bm{\theta}$.
@@
In the noise-dominated regime, the temporal trajectory samples a population of parameter locations around the path that the deterministic trajectory would have traced. That temporal wandering changes the search from single-value updates to quasi-Bayesian population-based method, in which the population distribution is shaped by the covariance of the noise process 49, 16.
@@
This algorithm combines a deterministic gradient step and a noise fluctuation 49. A simple stochastic differential equation describes the process
@@
The relative dominance of the deterministic gradient component compared with the noise component can be adjusted by changing $\eta$, the step-size weighting of the gradient term. As the $\eta$-weighted deterministic component declines, the method increasingly shifts from single-value updates to population-based updates. Here, dominance by temporal noise creates a similar effect to a spatially extended population 49, as described in the introduction to this section.
@@
The stochasticity from random sampling can help to escape local optima and find better parameter combinations 50. Thus, the method is often called stochastic gradient descent. In this process, a parameter update is
@@
in which $\nabla U$ is the true gradient, and $\bm{\xi}$ is the sampling noise of the gradient estimate. The variance of the sampling noise scales inversely with the batch sample size. Choosing a good batch size plays an important role in the optimization process 51, 52.
@@
As the batch size declines, the method increasingly shifts from deterministic single-value updates to stochastically sampled population-based updates. Here, dominance by temporal noise creates a similar effect to a spatially extended population 16, 53, as described in the introduction to this section.
@@
This section returns to population-based algorithms in relation to force and metric. I first introduce Gaussian processes, a population-based Bayesian method that weights alternative functions by how well they describe data rather than weighting alternative parameter vectors. The Bayesian weighting of alternatives provides a natural measure of uncertainty 54.
I then turn to the common Kalman filter method. That method is conceptually similar to Gaussian processes, using a time series sequence of learning updates to track a changing system rather than the typical one-shot learning update by which Gaussian processes describe static systems 55, 56.
@@
For a particular set of inputs, $\bm{\mathrm{X}}=\left\{\bm{\mathrm{x}}_{1},\dots,\bm{\mathrm{x}}_{N}\right\}$, the covariance matrix $\bm{\mathrm{K}}$ has entries $k(\bm{\mathrm{x}}_{i},\bm{\mathrm{x}}_{j})$. For these inputs, the associated mean values and covariance matrix define a particular multivariate Gaussian. That Gaussian is the Bayesian distribution over all candidate function values for these specific inputs.
@@
One typically uses a Gaussian process model to infer a static functional relation between inputs and outputs. A Kalman filter applies a similar approach to a dynamically changing system 57.
@@
The first Price equation term links the change in probability distribution, $\Delta\bm{\mathrm{q}}$, to the change in mean values, $\Delta_{\bm{\mathrm{f}}}\mskip 2.0mu\mskip 2.0mu\bar{\bm{\theta}}=\Delta\bm{\mathrm{q}}\cdot\bm{\theta}$. If we interpret the change in each frequency value, $q^{\prime}_{i}=q_{i}w_{i}$, as driven by relative performance, $w_{i}$, then can think of $\Delta\bm{\mathrm{q}}$ as the change in the probability distribution driven by the improvement caused by learning 24, 58, 59, 60.
@@
The variational Bayes method uses this approach 61. Let our candidate posterior probability distributions, $\hat{\bm{\mathrm{q}}}(\bm{\phi})$, be confined to a particular distributional form that depends on the parameters, $\bm{\phi}$. Then the problem becomes the search for $\bm{\phi}$ that minimizes the divergence of the assumed form for the posterior, $\hat{\bm{\mathrm{q}}}(\bm{\phi})$, from the true posterior for $\bm{\theta}$ given the data, $\bm{\mathrm{q}}^{\prime}(\bm{\theta}|\bm{\mathrm{D}})$.
We measure that difference between estimated and true posterior by the KL divergence, $\mathcal{D}\left(\hat{\bm{\mathrm{q}}}(\bm{\phi})||\bm{\mathrm{q}}^{\prime}\right)$, defined in eqn 13. This minimization problem is a variational method because we are minimizing the distance between the function $\hat{\bm{\mathrm{q}}}(\bm{\phi})$ and the true posterior, $\bm{\mathrm{q}}^{\prime}$.
Variational Bayes methods 61 provide a common approach to search for $\hat{\bm{\mathrm{q}}}(\bm{\phi})$. The details are simple but require a bit of notation to make the steps clear. In addition, we will use notation that can plug easily into the Price equation in the next section. Before starting, I list some notational shortcuts, in which $\bm{\mathrm{D}}$ refers to data
@@
${}\mskip 23.0muq(\bm{\theta}_{i},\bm{\mathrm{D}}){}$ $\mathrm{joint\ distribution\ \bm{\theta}\ and\ data}$
${}\mskip 23.0mup(\bm{\mathrm{D}}){}$ $\mathrm{probability\ of\ the\ data\ (constant)}$
@@
d’Alembert’s principle emphasizes how various forces in a conserved, static system balance, whereas most analyses focus on the dynamics of the moving parts. In other words, d’Alembert changes emphasis to the forces rather than the moving bodies 62. That perspective helps when the goal is to understand how a system works rather than in the calculation of the actual paths of motion.
@@
Friston built a unified brain theory on the principles of variational Bayes analysis. In essence, Friston sought a theory in which the minimization of a single quantity could unify three problems that had previously been treated as separate topics 63, 64, 36.
@@
Second, agents choose future actions that minimize expected free energy. To do that, they adopt an active inference policy that balances the tradeoff between exploitation and exploration 65. For exploitation, choose actions with likely outcomes that match current preferences. For exploration, choose actions that provide information to improve preferences, increasing the value of future outcomes.
@@
The exploitation component measures risk, the mismatch between preferred outcomes and actual outcomes. Friston measures risk by surprise, the mismatch between expected outcomes for a behavioral policy, $\pi$, and the actual outcomes. In probabilistic models, we write $\log q(o|C)$ for the surprise of an outcome, $o$, given an expected outcome, $C$. We get the total surprise by averaging over the frequency of the actual outcomes given the policy.
@@
In biology, Baldwin 66 was perhaps the first to recognize that such hierarchical separation can greatly accelerate the overall rate at which a system learns. Subsequent studies in biology have extended the idea that nonheritable developmental adjustments or learning by individuals explain many aspects of how populations have actually evolved over the history of life 67.
Alternatively, we may think of each individual or lower-level group as transmitting the parameter vector that it learns over time. Then the global process aggregates the learned parameter vectors from each lower-level unit, weighting each lower-level contribution by the relative performance associated with its transmitted vector. In biology, many studies of group selection emphasize the enhanced evolutionary potential that arises from hierarchical population structure 68, 69, 70, 71. Call this heritable learning.
@@
Many computational algorithms exploit the potential benefits of hierarchical learning. This section uses the Price equation and the FMB law to show how hierarchical learning fits into our simple and general framework for algorithmic learning and natural selection.
@@
Hinton & Nowlan 72 illustrated how individual learning transformed an unsolvable search into an easily solved one. Simplifying the original Hinton & Nowlan model, assume that each inherited genotype is a bit string of length 20. A particular target string is set as success. All other strings have the same low value of performance.
@@
Newer methods extend the Baldwin approach. For example, Fernando et al. 73 evolved a population of neural networks to solve a particular challenge. Each network inherits a parameter vector that sets the initial conditions, the way performance is calculated, and the learning rate. Each network then learns through a fixed number of standard stochastic gradient descent rounds of updates.
@@
These population models split into a nonheritable inner loop of learning within individuals and a heritable outer loop between individuals. The same split can be done with single-value parameter updates rather than populations. For example, model-agnostic meta-learning improves a single vector of initial weights for a learning process 74.
Those initial weights seed a task-specific inner loop with multiple steps of stochastic gradient descent. After evaluation of the final state, the updated parameter weights from this inner loop are discarded. Only the gradient of the inner-loop’s final performance with respect to the initial parameters is fed into the update process for the outer loop.
@@
The standard one-level FMB law holds in all of these cases. The force $\bm{\mathrm{f}}$ in the outer loop is the regression or gradient of fitness on the seed parameters. The inner-loop learning dynamics influences the calculation of fitness for a given seed parameter vector. Otherwise, the parameter updates from the inner loop are discarded.
The seed parameters of the outer loop often include hyperparameters, which are those parameters that control the inner-loop learning process rather than encode the parameters of the particular learned solution. Many other methods optimize hyperparameters and other attributes that control the architecture of the learning process in the outer loop and discard the parameter tuning in each inner-loop round 75, 76, 77, 78, 79, 80.
@@
To analyze multilevel learning, we recursively expand the Price equation to describe a population hierarchy 81, 82, 83. This subsection shows the steps.
@@
Group selection has been widely discussed in biology 68, 69, 70, 71, often using the Price equation’s recursive expansion 81, 82, 83. Many algorithmic learning methods have exploited the potential benefits of splitting populations into a multilevel hierarchy. This subsection mentions a few examples.
Commonly used evolutionary algorithms in machine learning can be extended by dividing populations into groups. These methods explicitly base their approach on the biological analogy of group selection or multilevel selection 84, 85.
Other population methods split learning into an outer loop and a heritable inner loop 86. The multistep inner loops may run within single agents of the population. Occasionally, the current state of one or more of the inner-loop agents is copied to seed some of the agents for a new population. For example, a currently strong agent may be cloned to replace a weaker agent. This approach combines the broad exploratory benefits of populations with the efficient improvement benefits of gradient-based approaches running within individual agents 87, 88.
@@
First, the look-ahead optimizer uses for its fast inner loop any common algorithm, such as stochastic gradient descent or Adam. Multiple inner-loop steps explore the local performance surface. The updated parameters then pass back to the slow outer loop, which adjusts the previously stored parameters in the direction of the inner-loop update. The next round of the fast inner loop begins with newly adjusted parameters, repeating the cycle 89.
@@
Second, we can interpret stochastic weight averaging within our hierarchical framework. The method first trains a model with stochastic gradient descent, forming the outer loop. Then it runs a further sequence of training steps in its inner loop. The outer loop is then updated to a final value by averaging over samples of the inner loop parameters 90.
@@
Third, deep Q networks optimize behavioral sequences. To value an action within the behavioral sequence of a focal network, the algorithm combines the observed reward for that action plus a predicted reward for future actions. To calculate the predicted future reward, the algorithm uses as a target the behavioral sequence of another network with a fixed parameter vector 91.

### v3

@@
Steven A. Frank**footnotemark: *
@@ Introduction
Learning algorithms pervade modern science. Machine learning improves neural networks through gradient descent. Evolution improves organisms through natural selection. Bayesian inference improves beliefs through probability updates. Despite decades of research in each field, the ultimate relations between these approaches remain unclear.
@@
These connections reflect a deeper geometric principle. Many learning algorithms face the same fundamental challenge: maximize improvement in performance minus a cost paid for distance moved in the parameter space. Here, we must account for two aspects of geometry. First, there may be a curved relation between parameters and performance. Second, constraints on movement in the parameter space induce a metric that alters how distance is measured. For example, lack of genetic variation in a particular direction constrains movement in that direction.
@@
The geometric structure of learning has been partially recognized in prior work. Fisher 9 and Rao 10 established that statistical parameter spaces can have intrinsic curvature. Amari 11 used this insight to develop natural gradient methods.
@@
This section’s alternative perspectives of dynamics adds a complementary axis to the local‑to‑population spectrum of methods introduced earlier. I develop this dynamics axis at the population scale, providing the most general conceptual frame. That population context prepares the ground for later discussion of particular algorithms, many of which blend local geometry with broader spatial or temporal scope.
@@
Equivalently, among all possible frequency changes, $\Delta\bm{\mathrm{q}}$, that produce the same improvement in performance, $\Delta_{\bm{\mathrm{f}}}\mskip 2.0mu\bar{w}$, the actual frequency changes for the given trait covariance matrix, $\bm{\mathrm{M}}$, lead to the shortest Fisher-Rao length. In biology, these optimality results are trait-based analogs of Fisher’s Fundamental Theorem for genetics 30, 31, 32, 33.
Of course, forces other than intrinsic performance can alter frequencies. The more we know about those other forces, such as environmental shifts or directional mutation in biology, the more accurately we can account for the consequences of those forces and improve inductive inference about the causal relation between parameters and performance 34, 35, 36.
@@
A particular algorithm can, of course, choose to modify how components of an update are calculated. However, populations set the foundation for analysis. Changes in parameter mean values, $\Delta\bar{\bm{\theta}}$, summarize updates. The metric $\bm{\mathrm{M}}$ typically arises from the covariance of alternative parameter values. Some methods use variational optimization and analogies with free energy, which links learning to various physical principles of dynamics 37, 38.
Second, many algorithms update a single parameter vector. These methods fill in the other end of the spectrum and its middle ground. The purely local strategy, such as Newton’s method, uses a metric and force gradient analyzed at a single point. Hybrid strategies choose metrics that incorporate a broader spatial scope, such as trust regions 39, 40 or the natural gradient 11.
@@
Finally, modern optimizers often include temporal scope. Extensions to stochastic gradient descent, such as Adam 8 and Nesterov 41, 42, explicitly use the history of updates to calculate a bias term, $\bm{\mathrm{b}}$. In these cases, bias explicitly applies physical notions of momentum, in which past movement of parameter values can push future updates beyond local traps on the performance surface.
@@
Covariance matrix algorithms (CMAs) often provide a good black-box optimization method 43. These evolution strategy (ES) methods proceed by analogy with biological evolution. One starts with a target parameter vector and then samples the performance values for a set of different parameter combinations around the target. That set forms a population from which one can calculate an updated target parameter by an empirically calculated covariance matrix and performance gradient, as in eqn 6.
These methods have three challenges: choosing sample points around the current target, estimating the performance gradient, and calculating the new target from the covariance matrix and performance gradient. I briefly summarize how the popular CMA-ES method handles these three challenges 44.
@@
The same $\bm{\mathrm{M}}\,\bm{\mathrm{f}}$ structure occurs in other evolution strategy (ES) learning algorithms, which include natural evolution strategies 45, large‑scale parallel ES 46, and separable low‑rank CMA‑ES 47. The next section moves to the local end of the spectrum, showing that the $\bm{\mathrm{M}}\,\bm{\mathrm{f}}$ structure remains when the population collapses to a single vector.
@@
The Broyden–Fletcher–Goldfarb–Shanno (BFGS) algorithm and its variants are widely used 48. After each step the method compares how the gradient changed to how the parameters moved. That comparison provides information about curvature.
@@
Newton’s local calculation of the curvature via the inverse Hessian is fragile. The local Hessian may not exist, it may change significantly over space, or it may not be invertible to provide the inverse curvature. Trust region methods may solve some of these problems by defining a local region of analysis 39, 40.
@@
The previous examples in this section match the $\bm{\mathrm{M}}\,\bm{\mathrm{f}}$ force-metric pattern in a simple and direct way. By contrast, the mirror descent algorithm has a more complicated geometry that does not obviously map directly to our $\bm{\mathrm{M}}\,\bm{\mathrm{f}}$ pattern 49, 50. However, the following derivations show that this algorithm does in fact use the same basic force-metric approach but with a more involved method to obtain the curvature metric.
@@
This update is a first-order approximation of the mirror descent method 49, 50. In effect, we are making a Newton-like step in an alternative mirror geometry with metric $\bm{\mathrm{H}}^{-1}_{\varphi}$, then mapping that step back to our original geometry for $\bm{\theta}$.
@@
In the noise-dominated regime, the temporal trajectory samples a population of parameter locations around the path that the deterministic trajectory would have traced. That temporal wandering changes the search from single-value updates to quasi-Bayesian population-based method, in which the population distribution is shaped by the covariance of the noise process 51, 16.
@@
This algorithm combines a deterministic gradient step and a noise fluctuation 51. A simple stochastic differential equation describes the process
@@
The relative dominance of the deterministic gradient component compared with the noise component can be adjusted by changing $\eta$, the step-size weighting of the gradient term. As the $\eta$-weighted deterministic component declines, the method increasingly shifts from single-value updates to population-based updates. Here, dominance by temporal noise creates a similar effect to a spatially extended population 51, as described in the introduction to this section.
@@
The stochasticity from random sampling can help to escape local optima and find better parameter combinations 52. Thus, the method is often called stochastic gradient descent. In this process, a parameter update is
@@
in which $\nabla U$ is the true gradient, and $\bm{\xi}$ is the sampling noise of the gradient estimate. The variance of the sampling noise scales inversely with the batch sample size. Choosing a good batch size plays an important role in the optimization process 53, 54.
@@
As the batch size declines, the method increasingly shifts from deterministic single-value updates to stochastically sampled population-based updates. Here, dominance by temporal noise creates a similar effect to a spatially extended population 16, 55, as described in the introduction to this section.
@@
This section returns to population-based algorithms in relation to force and metric. I first introduce Gaussian processes, a population-based Bayesian method that weights alternative functions by how well they describe data rather than weighting alternative parameter vectors. The Bayesian weighting of alternatives provides a natural measure of uncertainty 56.
I then turn to the common Kalman filter method. That method is conceptually similar to Gaussian processes, using a time series sequence of learning updates to track a changing system rather than the typical one-shot learning update by which Gaussian processes describe static systems 57, 58.
@@
We do not explicitly specify the functions, $\bm{\mathrm{g}}$, and their associated probabilities, $\bm{\mathrm{q}}$. Instead, we assume that, for an input vector $\bm{\mathrm{x}}$, the average output we expect over all functions is given by the mean curve $\mu(\bm{\mathrm{x}})=\operatorname{E}\left[g(\bm{\mathrm{x}})\right]$. The curve can be thought of as the plot of $\mu(\bm{\mathrm{x}})$ versus $\bm{\mathrm{x}}$.
@@
For inputs $\bm{\mathrm{X}}$, the associated vector of mean values and the covariance matrix define a particular multivariate Gaussian. That Gaussian is the Bayesian distribution over all candidate function values for these specific inputs.
@@
One typically uses a Gaussian process model to infer a static functional relation between inputs and outputs. A Kalman filter applies a similar approach to a dynamically changing system 59.
@@
The first Price equation term links the change in probability distribution, $\Delta\bm{\mathrm{q}}$, to the change in mean values, $\Delta_{\bm{\mathrm{f}}}\mskip 2.0mu\mskip 2.0mu\bar{\bm{\theta}}=\Delta\bm{\mathrm{q}}\cdot\bm{\theta}$. If we interpret the change in each frequency value, $q^{\prime}_{i}=q_{i}w_{i}$, as driven by relative performance, $w_{i}$, then can think of $\Delta\bm{\mathrm{q}}$ as the change in the probability distribution driven by the improvement caused by learning 24, 60, 61, 62, 63.
@@
The variational Bayes method uses this approach 64. Let our candidate posterior probability distributions, $\hat{\bm{\mathrm{q}}}(\bm{\phi})$, be confined to a particular distributional form that depends on the parameters, $\bm{\phi}$. Then the problem becomes the search for $\bm{\phi}$ that minimizes the divergence of the assumed form for the posterior, $\hat{\bm{\mathrm{q}}}(\bm{\phi})$, from the true posterior for $\bm{\theta}$ given the data, $\bm{\mathrm{q}}^{\prime}(\bm{\theta}|\bm{\mathrm{D}})$.
We measure that difference between estimated and true posterior by the KL divergence, $\mathcal{D}\left(\hat{\bm{\mathrm{q}}}(\bm{\phi})||\bm{\mathrm{q}}^{\prime}\right)$, defined in eqn 13. This minimization problem is a variational method because we are minimizing the divergence between the function $\hat{\bm{\mathrm{q}}}(\bm{\phi})$ and the true posterior, $\bm{\mathrm{q}}^{\prime}$.
Variational Bayes methods 64 provide a common approach to search for $\hat{\bm{\mathrm{q}}}(\bm{\phi})$. The details are simple but require a bit of notation to make the steps clear. In addition, we will use notation that can plug easily into the Price equation in the next section. Before starting, I list some notational shortcuts, in which $\bm{\mathrm{D}}$ refers to data
@@
${}\mskip 23.0muq(\bm{\theta}_{i},\bm{\mathrm{D}}){}$ $\mathrm{joint\ distribution}\ \bm{\theta}\ \mathrm{and\ data}$
${}\mskip 23.0mup(\bm{\mathrm{D}}){}$ $\mathrm{probability\ of\ the\ data}$
@@
d’Alembert’s principle emphasizes how various forces in a conserved, static system balance, whereas most analyses focus on the dynamics of the moving parts. In other words, d’Alembert changes emphasis to the forces rather than the moving bodies 65. That perspective helps when the goal is to understand how a system works rather than in the calculation of the actual paths of motion.
@@
Friston built a unified brain theory on the principles of variational Bayes analysis. In essence, Friston sought a theory in which the minimization of a single quantity could unify three problems that had previously been treated as separate topics 66, 67, 37.
@@
Second, agents choose future actions that minimize expected free energy. To do that, they adopt an active inference policy that balances the tradeoff between exploitation and exploration 68. For exploitation, choose actions with likely outcomes that match current preferences. For exploration, choose actions that provide information to improve preferences, increasing the value of future outcomes.
@@
The exploitation component measures risk, the mismatch between preferred outcomes and actual outcomes. Friston measures risk by surprise, the mismatch between expected outcomes for a behavioral policy, $\pi$, and the actual outcomes. In probabilistic models, we write $\log q(o|C)$ for the surprise of an outcome, $o$, given an expected outcome, $C$. We get the expected surprise by averaging over the frequency of the actual outcomes given the policy.
@@
In biology, Baldwin 69 was perhaps the first to recognize that such hierarchical separation can greatly accelerate the overall rate at which a system learns. Subsequent studies in biology have extended the idea that nonheritable developmental adjustments or learning by individuals explain many aspects of how populations have actually evolved over the history of life 70.
Alternatively, we may think of each individual or lower-level group as transmitting the parameter vector that it learns over time. Then the global process aggregates the learned parameter vectors from each lower-level unit, weighting each lower-level contribution by the relative performance associated with its transmitted vector. In biology, many studies of group selection emphasize the enhanced evolutionary potential that arises from hierarchical population structure 71, 72, 73, 74. Call this heritable learning.
@@
Many computational algorithms exploit the potential benefits of hierarchical learning 75. This section uses the Price equation and the FMB law to show how hierarchical learning fits into our simple and general framework for algorithmic learning and natural selection.
@@
Hinton & Nowlan 76 illustrated how individual learning transformed an unsolvable search into an easily solved one. Simplifying the original Hinton & Nowlan model, assume that each inherited genotype is a bit string of length 20. A particular target string is set as success. All other strings have the same low value of performance.
@@
Newer methods extend the Baldwin approach. For example, Fernando et al. 77 evolved a population of neural networks to solve a particular challenge. Each network inherits a parameter vector that sets the initial conditions, the way performance is calculated, and the learning rate. Each network then learns through a fixed number of standard stochastic gradient descent rounds of updates.
@@
These population models split into a nonheritable inner loop of learning within individuals and a heritable outer loop between individuals. The same split can be done with single-value parameter updates rather than populations. For example, model‑agnostic meta-learning improves a single vector of initial weights for a learning process 78.
Those initial weights seed a task‑specific inner loop with multiple steps of stochastic gradient descent. After evaluation of the final state, the updated parameter weights from this inner loop are discarded. Only the gradient of the inner-loop’s final performance with respect to the initial parameters is fed into the update process for the outer loop.
@@
The standard one‑level FMB law holds in all of these cases. The force $\bm{\mathrm{f}}$ in the outer loop is the regression or gradient of fitness on the seed parameters. The inner-loop learning dynamics influences the calculation of fitness for a given seed parameter vector. Otherwise, the parameter updates from the inner loop are discarded.
The seed parameters of the outer loop often include hyperparameters, which are those parameters that control the inner-loop learning process rather than encode the parameters of the particular learned solution. Many other methods optimize hyperparameters and other attributes that control the architecture of the learning process in the outer loop and discard the parameter tuning in each inner-loop round 79, 80, 81, 82, 83, 84.
@@
To analyze multilevel learning, we recursively expand the Price equation to describe a population hierarchy 85, 86, 87. This subsection shows the steps.
@@
Group selection has been widely discussed in biology 71, 72, 73, 74, often using the Price equation’s recursive expansion 85, 86, 87. Many algorithmic learning methods have exploited the potential benefits of splitting populations into a multilevel hierarchy. This subsection mentions a few examples.
Commonly used evolutionary algorithms in machine learning can be extended by dividing populations into groups. These methods explicitly base their approach on the biological analogy of group selection or multilevel selection 88, 89.
Other population methods split learning into an outer loop and a heritable inner loop 90. The multistep inner loops may run within single agents of the population. Occasionally, the current state of one or more of the inner-loop agents is copied to seed some of the agents for a new population. For example, a currently strong agent may be cloned to replace a weaker agent. This approach combines the broad exploratory benefits of populations with the efficient improvement benefits of gradient-based approaches running within individual agents 91, 92.
@@
First, the look-ahead optimizer uses for its fast inner loop any common algorithm, such as stochastic gradient descent or Adam. Multiple inner-loop steps explore the local performance surface. The updated parameters then pass back to the slow outer loop, which adjusts the previously stored parameters in the direction of the inner-loop update. The next round of the fast inner loop begins with newly adjusted parameters, repeating the cycle 93.
@@
Second, we can interpret stochastic weight averaging within our hierarchical framework. The method first trains a model with stochastic gradient descent, forming the outer loop. Then it runs a further sequence of training steps in its inner loop. The outer loop is then updated to a final value by averaging over samples of the inner loop parameters 94.
@@
Third, deep Q networks optimize behavioral sequences. To value an action within the behavioral sequence of a focal network, the algorithm combines the observed reward for that action plus a predicted reward for future actions. To calculate the predicted future reward, the algorithm uses as a target the behavioral sequence of another network with a fixed parameter vector 95.